# Semiclassical Backreaction: A Qualitative Assessment

### v1

@@ I Introduction
The latter example is particularly important, since it suggests that black holes radiate thermally and thus, by energy conservation, lose mass and evaporate. Such a situation lies precisely outside the regime of validity of the semiclassical regime of quantum mechanics described above since the classical background (the metric around a Schwarzschild black hole for instance) must be sensitive to the production of Hawking quanta. This backreaction of quantum radiation on the classical background that has generated it lacks a satisfying description since a full theory of quantum gravity is unavailable for the time being. One possible way of taking quantum backreaction into account is to replace the energy-momentum tensor on the right hand side of the Einstein equations with its quantum expectation value taken in a properly chosen quantum state. This of course comes with its own difficulties because surely such a mean field approximation would only be valid as long as the quantum state stays sufficiently regular and doesn’t develop a spatially multi-modal character (the analog of a double-humped spatial probability density in quantum mechanics). [^2] Not to mention the many subtleties relating to the renormalization of the usual quantum field theoretic divergences accompanying correlation functions in the coincident limit [15].
@@
Of course there are various other ways of taking backreaction into account, with the help of Langevin or Fokker-Planck equations [22, 23, 24], by integrating out quantum fluctuations and using an effective action to determine the background dynamics [25, 26], or by using open quantum field theory methods akin to the so-called in-in or Schwinger-Kheldysh formalism [27, 23, 28]. However all of these methods are quite unwieldy and often times lead to computations that are impossible to perform in practice. Moreover, to our knowledge none of the above semiclassical backreaction methods has been checked against a fully quantum mechanical treatment of the coupled system (either because of the intrinsic complexity of the task or because such a treatment is unavailable as in the case of gravity).
@@ II Motivation
For any finite value of $\hbar$ we expect a finite value of the quantum break time, which was argued to scale as $t_{q}\sim 1/(\omega\alpha^{2}\mathcal{N})$ in [29]. Since a full quantum treatment is usually impossible, simply because of the dimensionality of the problem, various semiclassical approximations have been proposed to study quantum corrections in the limit of large occupation number $\mathcal{N}$ and be able to evolve the problem beyond the quantum break time. Generically, all of these methods assume $\phi$ to be classical and $\chi$ to be quantum, and prescribe the way in which these two fundamentally different degrees of freedom interact with one another. However, since in most cases we don’t have an exact quantum field theoretic computation to compare to, the validity of these approximations is untested. As mentioned in the introduction, we are interested in comparing two such popular methods to a full quantum mechanical calculation. We are thus forced to limit ourselves to a toy model of two coupled quantum harmonic oscillators. This can be thought of as the field theory model of Eq. (1) restricted to the homogeneous mode of the field $\phi$ and just one of the real modes of the field $\chi$. We are thus interested in computing the effect that the production of quanta in a $\chi$ field mode has on the dynamics of the homogeneous mode of the classical field $\phi$. In other words we are interested in the semiclassical backreaction of $\chi$ on $\phi$ in a mini-superspace approximation. In the next section, we introduce the toy model that we will consider as well as the two semiclassical approximation schemes. From now on we will focus exclusively on this simplified system but we will also be touching upon concepts that are important to the more general case outlined here in this section.
@@ III Model
Let us pause for a moment to define what we mean by entanglement here and more generally throughout this paper. Generally, an entangled quantum state is defined in opposition to a product (or separable) state. Thus, in the case of our toy model, an entangled state will be a state |$\ket{\psi}$ that cannot be written as a tensor product |$\ket{\rho}_{x}\otimes\ket{\sigma}_{y}$ of an $x$ state and a $y$ state. There are many measures of entanglement including purity, entanglement entropy, and more generally Renyi entropies [37], but here we will adopt a less sophisticated approach. We will call a state that is separable up to terms of order $\lambda$ a weakly entangled state, and conversely, a state that does not satisfy this property will be called strongly entangled.
@@ III.1 Numerical methods
In what follows we will compare the fully quantum mechanical evolution of $\langle\hat{x}(t)\rangle$ obtained from a numerical solution of the Schrödinger equation with Hamiltonian (5), to the one predicted by two semiclassical approximation schemes for including the effects of the quantum backreaction of $y$ on the (classical) dynamics of $\langle\hat{x}(t)\rangle$. We will refer to these two methods as the Mean Field and Truncated Wigner methods. [^5] Note that what call MF method is known in the litterature by many names: Hartree, Hartree-Fock or simply semiclassical approximation (especially in a semiclassical gravity context). Likewise the TW method is sometimes known as the stochastic or statistical method.
@@ III.1.1 Quantum Mechanics
Starting from these initial conditions we will solve the Schrödinger equation numerically using a second-order symplectic integrator (see Appendix B). At any point in time we can then output $\langle\hat{x}(t)\rangle$ and use this to assess the performance of the semiclassical approximation schemes. We outline these schemes next.
@@ III.1.2 Semiclassical approximations
    At this point $x(t)$ can be chosen to be $x_{0}\cos t$ but in fact it can be any arbitrary background function of time such that $x(0)=x_{0}$. The appropriate initial conditions for $z(t)$, corresponding to the choice of the adiabatic ground state for $y$, are $z(0)=-i\left(\left(\Omega/\omega\right)^{2}+\lambda x_{0}^{2}\right)^{-1/4}$ and $\dot{z}(0)=\left(\left(\Omega/\omega\right)^{2}+\lambda x_{0}^{2}\right)^{1/4}$. As long as the function $x(t)$ is fixed, the solution $z(t)$ of Eq. (9) is enough to specify the exact quantum dynamics of $y$ i.e. to exactly compute all of its correlation functions. For instance $\langle\hat{y}(t)^{2}\rangle=|z(t)|^{2}$. However, in this case we are precisely interested in how the background $x$ is corrected, or backreacted upon, by the quantum excitations of $y$. Within the Mean Field (MF) approximation, we include this correction in the classical equation of motion for $x$, Eq. (7), by replacing $y^{2}$ with $\langle\hat{y}^{2}\rangle=|z|^{2}$. The solution, $x(t)$, of the resulting coupled system of equations will thus deviate from the purely classical harmonic and will also appear in Eq. (9). In summary, the set of equations to be solved for in the MF approximation will be
@@
    and can be interpreted as a joint probability distribution for the initial phase space variables of our system of interest. The idea of the Truncated Wigner (TW) method is to approximate the time-evolved Wigner function by a truncated sum of $\delta$ functions,
@@ IV Results
We observe that although the heuristic linear scaling is sometimes correct, the influence of classical instabilities can be considerable, spoiling this prediction. In particular, as has been noted before, classically unstable systems have break times that scale with $\log N$ instead [45, 33].
@@ V Summary and Discussion
The previous observations have important consequences for some of the major domains of application of these two semiclassical methods. Taking the example of (p)reheating with which we started our discussion in Sec. II, we are now able to reinterpret the standard methodology in terms of our analysis. In fact, what is typically done in the literature of preheating, is to compute particle production from parametric resonance of classical fluctuations (typically obtained through cosmological lattice simulations), averaged over a variety of initializations [21]. In essence this is an application of the Truncated Wigner method to cosmological quantum fluctuations. We now understand why this is the correct choice as parametric resonance and classical instability are central to preheating, and one can expect a fair amount of entanglement between the inflaton and the resonant vacuum modes. The stable analogue of preheating is reheating [54]. Our expectation is that a mean-field approach should be appropriate in this scenario. Note that the period of reheating has traditionally been studied by assuming cubic couplings $\propto\phi\chi^{2}$, instead of the quartic coupling under consideration here. It is an interesting question to understand how reheating calculations would be altered by using the MF approach. We leave this interesting question for future work. Other scenarios that have been treated using either of the approaches in this paper include the quantum decay of breather states (MF method) [41], kink-antikink scatterings in a quantum vacuum (MF method) [40], and false vacuum decay (TW method) [20]. It would be interesting to revisit some of these cases and assess their validity given the results in this work. Particularly worrisome is that many breather states under consideration in the literature have nonperturbative collective couplings where $\lambda N=\mathcal{O}(1)$.
@@
One of the main limitations of this work is its reliance on the assumption that $\langle\hat{x}(t)\rangle$ computed in a full quantum treatment, should always be a good benchmark against which to compare the performance of different semiclassical approximation methods. This is in fact only true as long the wavefunction $\psi(x,y)$ for the coupled system is well localized and more or less monomodal in the $x$ direction. When this ceases to be true, it is clear that the average expectation value cannot possibly represent the dynamics of a classical variable. This introduces a more fundamental upper limit for the duration of validity of any semiclassical approximation, irrespectively of the associated value of $\delta(t)$ (see Eq. (16)). While we have checked that the wavefunction remains approximately Gaussian in the $x$ direction for the duration of our simulation for some parameter choices, we leave a more exhaustive and rigorous study of this issue for future work.
@@ Acknowledgments
We would like to thank Eugenio Bianchi, Martin Bojowald, Tanmay Vachaspati, and Juan Sebastien Valbuena for useful discussions. F. D. acknowledges the support from the Departament de Recerca i Universitats from Generalitat de Catalunya to the Grup de Recerca 00649 (Codi: 2021 SGR 00649) and funding from the ESF under the program Ayudas predoctorales of the Ministerio de Ciencia e Innovación PRE2020-094420. G. Z. is supported by SONATA BIS Grant No. 2023/50/E/ST2/00231 from the Polish National Science Centre.
@@ A.3 Trunctated Wigner
Appendix B Numerics and performance tests

### v2

@@ I Introduction
The latter example is particularly important, since it suggests that black holes radiate thermally and thus, by energy conservation, lose mass and evaporate. Such a situation lies precisely outside the regime of validity of the semiclassical regime of quantum mechanics described above since the classical background (the metric around a Schwarzschild black hole for instance) must be sensitive to the production of Hawking quanta. This backreaction of quantum radiation on the classical background that has generated it, lacks a satisfying description since a full theory of quantum gravity is unavailable for the time being. One possible way of taking quantum backreaction into account is to replace the energy-momentum tensor on the right hand side of the Einstein equations with its quantum expectation value taken in a properly chosen quantum state. This of course comes with its own difficulties because surely such a mean field approximation would only be valid as long as the quantum state stays sufficiently regular and doesn’t develop a spatially multi-modal character (the analog of a double-humped spatial probability density in quantum mechanics). [^2] Not to mention the many subtleties relating to the renormalization of the usual quantum field theoretic divergences accompanying correlation functions in the coincident limit [15].
@@
Of course there are various other ways of taking backreaction into account, with the help of Langevin or Fokker-Planck equations [22, 23, 24], by integrating out quantum fluctuations and using an effective action to determine the background dynamics [25, 26], or by using open quantum field theory methods akin to the so-called in-in or Schwinger-Kheldysh formalism [27, 23, 28]. However all of these methods are quite unwieldy and often times lead to computations that, although in principle doable to arbitrary precision (see for instance Ref. [29] where the reflection coefficient on a so-called reflectionless kink is computed in a fully quantum field theoretic manner and backreaction is included), are difficult to perform in practice. Moreover, to our knowledge none of the above semiclassical backreaction methods has been checked against a fully quantum mechanical treatment of the coupled system (either because of the intrinsic complexity of the task or because such a treatment is unavailable as in the case of gravity).
@@ II Motivation
For any finite value of $\hbar$ we expect a finite value of the quantum break time, which was argued to scale as $t_{q}\sim 1/(\omega\alpha^{2}\mathcal{N})$ in [30]. Since a full quantum treatment is usually difficult, [^5] Although some attempts have been made to study this problem by explicitly constructing high occupancy coherent states in an interacting quantum field theoretic setting [38, 35, 39, 40, 41]. simply because of the dimensionality of the problem, various semiclassical approximations have been proposed to study quantum corrections in the limit of large occupation number $\mathcal{N}$ and be able to evolve the problem beyond the quantum break time. Generically, all of these methods assume $\phi$ to be classical and $\chi$ to be quantum, and prescribe the way in which these two fundamentally different degrees of freedom interact with one another. However, since in most cases we don’t have an exact quantum field theoretic computation to compare to, the validity of these approximations is untested. As mentioned in the introduction, we are interested in comparing two such popular methods to a full quantum mechanical calculation. We are thus forced to limit ourselves to a toy model of two coupled quantum harmonic oscillators. This can be thought of as the field theory model of Eq. (1) restricted to the homogeneous mode of the field $\phi$ and just one of the real modes of the field $\chi$. We are thus interested in computing the effect that the production of quanta in a $\chi$ field mode has on the dynamics of the homogeneous mode of the classical field $\phi$. In other words we are interested in the semiclassical backreaction of $\chi$ on $\phi$ in a minisuperspace approximation. In the next section, we introduce the toy model that we will consider as well as the two semiclassical approximation schemes. From now on we will focus exclusively on this simplified system but we will also be touching upon concepts that are important to the more general case outlined here in this section.
@@ III Model
Let us pause for a moment to define what we mean by entanglement here and more generally throughout this paper. Generally, an entangled quantum state is defined in opposition to a product (or separable) state. Thus, in the case of our toy model, an entangled state will be a state |$\ket{\psi}$ that cannot be written as a tensor product |$\ket{\rho}_{x}\otimes\ket{\sigma}_{y}$ of an $x$ state and a $y$ state. There are many measures of entanglement including purity, entanglement entropy, and more generally Renyi entropies [42], but here we will mostly adopt a less sophisticated approach. We will call a state that is separable up to terms of order $\lambda$ a weakly entangled state, and conversely, a state that does not satisfy this property will be called strongly entangled. At the end of Sec. IV we will check that this naive approach agrees with more quantitative measures of quantum entanglement.
@@ III.1 Numerical methods
In what follows we will compare the fully quantum mechanical evolution of $\langle\hat{x}(t)\rangle$ obtained from a numerical solution of the Schrödinger equation with Hamiltonian (5), to the one predicted by two semiclassical approximation schemes for including the effects of the quantum backreaction of $y$ on the (classical) dynamics of $\langle\hat{x}(t)\rangle$. We will refer to these two methods as the Mean Field (MF) and Truncated Wigner (TW) methods. [^6] Note that what we call MF method is known in the literature by many names: Hartree, Hartree-Fock or simply semiclassical approximation (especially in a semiclassical gravity context). Likewise the TW method is sometimes known as the stochastic or statistical method.
@@ III.1.1 Quantum Mechanics
Starting from these initial conditions we will solve the Schrödinger equation numerically using a second-order symplectic integrator (see Appendix C). At any point in time we can then output $\langle\hat{x}(t)\rangle$ and use this to assess the performance of the semiclassical approximation schemes. We outline these schemes next.
@@ III.1.2 Semiclassical approximations
    At this point $x(t)$ can be chosen to be $x_{0}\cos t$ but in fact it can be any arbitrary background function of time such that $x(0)=x_{0}$. The appropriate initial conditions for $z(t)$, corresponding to the choice of the adiabatic ground state for $y$, are $z(0)=-i\left(\left(\Omega/\omega\right)^{2}+\lambda x_{0}^{2}\right)^{-1/4}$ and $\dot{z}(0)=\left(\left(\Omega/\omega\right)^{2}+\lambda x_{0}^{2}\right)^{1/4}$. As long as the function $x(t)$ is fixed, the solution $z(t)$ of Eq. (9) is enough to specify the exact quantum dynamics of $y$ i.e. to exactly compute all of its correlation functions. For instance $\langle\hat{y}(t)^{2}\rangle=|z(t)|^{2}$. However, in this case we are precisely interested in how the background $x$ is corrected, or backreacted upon, by the quantum excitations of $y$. Within the MF approximation, we include this correction in the classical equation of motion for $x$, Eq. (7), by replacing $y^{2}$ with $\langle\hat{y}^{2}\rangle=|z|^{2}$. The solution, $x(t)$, of the resulting coupled system of equations will thus deviate from the purely classical harmonic and will also appear in Eq. (9). In summary, the set of equations to be solved for in the MF approximation will be
@@
    and can be interpreted as a joint probability distribution for the initial phase space variables of our system of interest. The idea of the TW method is to approximate the time-evolved Wigner function by a truncated sum of $\delta$ functions,
@@ IV Results
To summarize our results and to make the connection with quantum entanglement a bit more precise, in Table 1 we give the break times as well as the entanglement entropy (at some fiducial time) for a few prototypical values of the parameters. The entanglement entropy of the system is computed as the Von Neumann entropy of the reduced density matrix, $\hat{\rho}_{x}$,
$\hat{\rho}_{x}=\sum_{m}\bra{m}\hat{\rho}\ket{m},$ (19)
where |$\ket{m}$ are the eigenstates of the $y$ quantum harmonic oscillator and $\hat{\rho}$ is the full density matrix of the coupled system. As the Hilbert space has infinite dimension we limit the computation to a maximum quantum number $m_{max}=28$, making sure that the state is contained up to a few parts in $10^{2}$ in this reduced space. Then the entanglement entropy is given by
$S_{e}=-\sum_{i}\lambda_{i}\log(\lambda_{i}),$ (20)
where $\lambda_{i}$ are the eigenvalues of $\hat{\rho}_{x}$. Heuristically, $S_{e}$ measures the uncertainty in the “classical” $x$ degree of freedom due to the interaction with the “quantum” $y$ degree of freedom. Its maximum value is related to the dimensionality of the Hilbert space: in our case $S_{e,max}=\ln(m_{max})=3.33$. If and only if there is no entanglement (so that the state is separable) is it exactly 00 0. Note that this way of estimating quantum entanglement is just one of many possibilities [42].
$\lambda$ $x_{0}$ $\Omega/\omega$ Stability $t_{q}$ $t_{MF}$ $t_{TW}$ $S_{e}$ at $t=25$
0.1 5 0.4 Unstable 6 9 17 1.86
0.05 8 0.4 Unstable 9 12 22 1.64
0.03 5 0.4 Stable 14 83 73 0.31
0.1 5 0.8 Unstable 10 16 25 1.21
0.04 6 0.8 Unstable 13 20 32 1.47
0.05 8 0.8 Stable 20 127 105 0.70
0.03 5 1.2 Stable 40 3304 164 0.026
0.04 6 1.2 Stable 31 1385 140 0.068
0.08 10 1.2 Unstable 14 22 34 0.13
Table 1: A summary of some prototypical results from our numerical computations. The influence of stability on entanglement, and in turn the influence of entanglement on the break times, is evident.
The results in Table 1 contain the most salient features of our results. Namely, unstable modes result in larger entanglement, which in turn reduces the break times of the various methods. As suggested earlier, when entanglement is small, one typically has $t_{MF}>t_{TW}$. The quantum break time $t_{q}$ is always smaller than both $t_{MF}$ and $t_{TW}$, showing that the methods give at least some improvement with respect to a purely classical treatment. Note in particular that for the stable cases of $\Omega/\omega=1.2$, quantum entanglement is extremely small and $t_{MF}$ is orders of magnitude larger than both $t_{q}$ and $t_{TW}$.
@@
We observe that although the heuristic linear scaling is sometimes correct, the influence of classical instabilities can be considerable, spoiling this prediction. In particular, as has been noted before, classically unstable systems have break times that scale with $\log N$ instead [50, 34]. For certain values of the collective coupling, one can have a situation where the system exits an instability band during the double scaling limit. There, we observe a transition between the two types of dependence ($\propto N$ and $\propto\log N$), e.g. the rightmost plot of Fig. 7.
@@ V Summary and Discussion
The previous observations have important consequences for some of the major domains of application of these two semiclassical methods. Taking the example of (p)reheating with which we started our discussion in Sec. II, we are now able to reinterpret the standard methodology in terms of our analysis. In fact, what is typically done in the literature of preheating, is to compute particle production from parametric resonance of classical fluctuations (typically obtained through cosmological lattice simulations), averaged over a variety of initializations [21]. In essence this is an application of the Truncated Wigner method to cosmological quantum fluctuations. We now understand why this is the correct choice as parametric resonance and classical instability are central to preheating, and one can expect a fair amount of entanglement between the inflaton and the resonant vacuum modes. The stable analogue of preheating is reheating [59]. Our expectation is that a mean-field approach should be appropriate in this scenario. Note that the period of reheating has traditionally been studied by assuming cubic couplings $\propto\phi\chi^{2}$, instead of the quartic coupling under consideration here. It is an interesting question to understand how reheating calculations would be altered by using the MF approach. We leave this interesting question for future work.
Other scenarios that have been treated using either of the approaches in this paper include the quantum decay of breather states (MF method) [46], kink-antikink scatterings in a quantum vacuum (MF method) [45], and false vacuum decay (TW method) [20]. It would be interesting to revisit some of these cases and assess their validity given the results in this work. [^9] Many breather states under consideration in the literature have nonperturbative collective couplings where $\lambda N=\mathcal{O}(1)$. It has been suggested that the threshold $\lambda N=\mathcal{O}(1)$ characterizes a class of macroscopic objects that saturate the entropy bound coming from $2\to N$ scattering amplitudes of the theory under consideration [60]. It is also important to mention that the quantum stability of soliton, oscillon, and breather backgrounds has been thoroughly investigated in the literature using perturbative methods as for instance in Ref. [61], and more recently in Refs. [62, 63] where it was argued that semiclassical methods may obscure important details about the quantum state of the background. For instance it was found that the quantum breather decay discussed using semiclassical MF methods in Ref. [46], resulted from an implicitly non-coherent choice of state for the background, rather than from an intrinsic quantum instability of the breather.
@@
One of the main limitations of this work is its reliance on the assumption that $\langle\hat{x}(t)\rangle$ computed in a full quantum treatment, should always be a good benchmark against which to compare the performance of different semiclassical approximation methods. This is in fact only true as long the wavefunction $\psi(x,y)$ for the coupled system is well localized and more or less monomodal in the $x$ direction. When this ceases to be true, it is clear that the average expectation value cannot possibly represent the dynamics of a classical variable. This introduces a more fundamental upper limit for the duration of validity of any semiclassical approximation, irrespectively of the associated value of $\delta(t)$ (see Eq. (16)). While we have checked that the wavefunction remains approximately Gaussian in the $x$ direction for the duration of our simulation for some parameter choices, we leave a more exhaustive and rigorous study of this issue for future work. Another important limitation is the fact that it is unclear how to extend our results to the many-mode case and eventually to field theory. Appendix B gives some partial answers in the case where the “classical” mode is coupled to not one but two “quantum” modes. Again the presence of unstable modes seems to shorten the validity of the semiclassical methods under consideration but more work is needed in order to really understand the net effect of stable and unstable modes on the background in a many mode scenario. We also plan to tackle this problem in the near future. The results of Appendix B do provide some indication that classical instability and entanglement continue to play an important role in field theory, as it does in the quantum mechanical system under consideration here.
@@ Acknowledgments
We would like to thank Eugenio Bianchi, Martin Bojowald, Tanmay Vachaspati, and Juan Sebastien Valbuena for useful discussions. F. D. acknowledges the support from the Departament de Recerca i Universitats from Generalitat de Catalunya to the Grup de Recerca 00649 (Codi: 2021 SGR 00649) and funding from the ESF under the program Ayudas predoctorales of the Ministerio de Ciencia e Innovación PRE2020-094420. G. Z. is supported by SONATA BIS Grant No. 2023/50/E/ST2/00231 from the Polish National Science Centre. This work is part of the doctoral thesis of F. D. within the framework of the Doctoral Program in Physics at the Autonomous University of Barcelona. The data used to produce the figures in this work are available at [65].
@@ A.3 Trunctated Wigner
Appendix B Systems with more degrees of freedom
This appendix constitutes a preliminary investigation of how our results extend to systems with more degrees of freedom. In field theory, there are many modes, of which some can be stable, while others can be unstable. To understand how our results extend to quantum field theory it is thus important to introduce more degrees of freedom in our analysis.
As a first step, we analyze a three mode system defined through the dimensionless Hamiltonian
$H=-\frac{1}{2}\partial^{2}_{x}+\frac{1}{2}x^{2}-\frac{1}{2}\partial^{2}_{y}+\frac{1}{2}\left[\left(\Omega_{1}/\omega\right)^{2}+\lambda x^{2}\right]y^{2}-\frac{1}{2}\partial^{2}_{z}+\frac{1}{2}\left[\left(\Omega_{2}/\omega\right)^{2}+\lambda x^{2}\right]z^{2}.$ (47)
Here both the $y$ and $z$ degree of freedom should be interpreted as quantum modes, while $x$ still corresponds to the classical background. We perform the computations in the same way as for the two-mode system, focusing on the break times of the different methods as well as on the entanglement entropy of $x$ (this time the density matrix is partially traced over both the $y$ and $z$ degrees of freedom whose Hilbert space has been restricted to a finite dimensional subspace). We are particularly interested in cases where: (i) both modes are classically stable, (ii) both modes are classically unstable, and (iii) one mode is unstable while the other is stable. Results are shown in Table 2.
$\lambda$ $x_{0}$ $\Omega_{1}/\omega$ $\Omega_{2}/\omega$ Stability $t_{q}$ $t_{MF}$ $t_{TW}$ $S_{e}$ at $t=25$
0.1 5 0.4 0.8 UU 4 9 17 2.93
0.05 8 0.4 0.8 US 7 12 25 2.37
0.04 6 0.8 1.2 US 11 19 31 1.62
0.03 5 0.4 1.2 SS 11 83 70 0.34
Table 2: Some prototypical results for the three mode system. Three situations can be distinguished: UU, when both modes are unstable, US, when the first mode is unstable while the second one is stable, and SS when both modes are stable. We can readily compare these results to those of Table 1, for the two mode system. Again the influence of the existence of unstable modes on the break times is evident.
Our conclusions about the influence of classical instability and entanglement on the break times seem to extend to the three mode system. Comparing the above results to the analogue ones for the two mode case (see Table 1), we see that the “most entangling” mode ultimately determines the entanglement entropy and break times of the three mode system. However the amount of unstable modes also seems to play a role. It is in particular unclear to what extent the presence of more stable modes can wash out the effect of the instability and extend the validity of semiclassical methods. Unfortunately incorporating more modes quickly becomes computationally expensive, and is thus beyond the scope of this work.
Appendix C Numerics and performance tests